### v1

# High-Performance Wavelength Division Multiplexers Enabled by Co-Optimized Inverse Design

Abstract
Wavelength division multiplexers are fundamental to the functioning and performance of integrated photonic circuits, with applications ranging from optical interconnects to sensing and quantum technologies. Current solutions are limited by trade-offs between channel spacing, crosstalk, insertion loss, and device footprint. Here, we develop a novel design approach that co-optimizes inverse-designed wavelength division multiplexers and distributed Bragg gratings to achieve ultra-low crosstalk without compromising insertion loss. We experimentally demonstrate less than -40 dB crosstalk for wavelength channel spacing of 15 nm in foundry-compatible silicon devices across the telecommunications C- and L-bands. Our design process is highly adaptable, allowing for seamless scaling to a greater number of output channels, different spectral windows, and easy translation across various material platforms.
1 Introduction
Integrated photonics has emerged as a scalable platform for data communication, becoming prevalent for use in optical interconnects for data centers [1]. To scale transmission bandwidth, many systems have exploited multi-wavelength data carriers—such as frequency comb or quantum dot mode-locked lasers—to achieve narrowly spaced wavelength channels spanning up to an octave, enabling hundreds of multiplexed information channels [2, 3, 4, 5]. Despite the benefits that a dense frequency source offers for optical data communications, the presence of widely spanned narrow channels introduces increased complexity in preserving the signal integrity required for on-chip modulation and data encoding. There have been many approaches to on-chip wavelength de-multiplexing (WDM) including but not limited to arrayed waveguide gratings [6], thermally tuned ring resonators [7, 8] and inverse design [9, 10]. To preserve signal integrity, it is essential for WDM solutions to exhibit low insertion loss and crosstalk. Arrayed waveguide gratings (AWGs) operate on principles from free space diffraction gratings and have been used in silicon integrated photonic circuits for many decades [11]. More recent advances in different material platforms have motivated the development of AWGs with materials such as silicon nitride [12] and thin film lithium niobate [13] which have low loss and nonlinearity, useful for other optical interconnect components. Although AWGs can achieve sub-3 dB insertion loss [6, 13] and less than 4 nm channel spacing [13], the requirement for large propagation lengths in both the arrayed waveguides and the free propagation region restricts the miniaturization of AWGs. This results in AWG solutions with device footprints on the order of 100 100 m2. Tunable ring resonators can act as filterbanks, using cascaded rings to achieve flat-top responses for pass band channels [7, 8, 14]. In this approach, the refractive index of the rings are thermo-optically tuned to align the resonances with the input channels. Insertion loss of 1.5 dB, extinction ratio greater than 35 dB and channel spacing of 124 GHz have been demonstrated [8]. Although the singular rings have small radii (around 6-12 m [7, 8]), this approach requires extra energy for the thermal tuning and two rings per output port, increasing the footprint.
Among existing state-of-the-art solutions for wavelength de-multiplexing is inverse design. Advancing technology has motivated higher data rates in optical interconnects and consequently more creative design approaches have materialized. Inverse design allows researchers to explore the entire design space, creating compact and efficient photonic devices with optimized dielectric distributions [15]. Using the adjoint method (which only requires two simulations per design iteration) and accelerated electromagnetic simulations, photonic devices can be optimized for a variety of functionalities [15]. Devices demonstrated include grating couplers [16], cavities [17], filters [18], spatial mode de-multiplexers [19] and wavelength de-multiplexers [9, 10]. Despite the irregular geometries that arise from inverse design, the scalability of the resulting devices has been proven through the enforcement of foundry fabrication constraints during optimization [20, 21, 22]. Although this approach can outperform classical designs and overcome trade-offs between metrics such as crosstalk, channel spacing, insertion loss, and device area, the best inverse design WDM to date is limited in channel spacing to 20 nm and crosstalk at -26 dB [23].
Recent advancements in simulation capacity, specifically GPU-accelerated large-scale FDTD (Finite Difference Time Domain) solvers, have opened the door for the next generation of photonic devices. In free space optics, the design of large-area metasurfaces was previously bottlenecked by simulation volume and computation time. Creative solutions for simulating wide-area metasurfaces [24, 25] and GPU-optimized FDTD solvers [26, 27] have enabled significant progress in free-space optics, which greatly benefits from these larger metasurfaces. In this work, we demonstrate the benefit this recently expanded simulation capacity can have on integrated photonics: multifunctional optimization. Our wavelength de-multiplexer is an example of the highly performing devices that can result from adding a component with some existing functionality to the large simulation region. The inclusion of other photonic components as a part of the inverse design process and optimizing with respect to this existing functionality helps to enable the next generation of photonic devices.
Distributed Bragg waveguide gratings form a one-dimension photonic band gap through periodic perturbations of the refractive index. This band gap makes Bragg gratings excellent candidates for wavelength selective filtering, with near unity transmission of the very narrow pass band (can be as small as 0.5 nm [28]). Bragg filters have yet to be fully explored in WDM systems due to the undesirable, also near unity, reflection of the wavelengths in the rejected band. Here, we introduce a new approach to the inverse design of WDMs with ultra-low crosstalk through co-optimization with Bragg filters. By initializing the design region with the Bragg filters at the WDM output, the reflection from the Bragg gratings is included in the loss function used for optimization and we achieve narrowband WDM performance by using inverse design to efficiently couple to the Bragg modes. This co-design principle is illustrated in Figure 1 and 2a.
2 Design Approach
Prior to inverse design, the Bragg gratings are designed to align with the desired WDM output channels. Equations 1 and 2 govern the bandwidth and central wavelength of the Bragg gratings, given the effective refractive indices neff1 and neff2 of the unperturbed and perturbed waveguide sections, the periodicity of the perturbations, , and the length of the grating . Here, represents the average effective index defined by and is the coupling coefficient equal to .
For our demonstration, we design a 2 channel WDM with a channel spacing of 15 nm. Given this channel separation, the two Bragg gratings are designed to have central wavelengths that are 15 nm apart. The bandwidths of the grating are designed such that the edge of one filter aligns with the edge of the other, which gives a bandwidth of 15 nm. The input waveguide width is 500 nm, for single mode operation at telecom wavelengths. The Bragg gratings have widths of 1 m with perturbations of 500 nm. The performance and spectral alignment of the filters are confirmed using FDTD simulations (Spins Photonics fdtdz) and imported into the optimization as an initial condition. The design region is defined as a 12 m 12 m region with the two filters at the output ports. The number of Bragg periods, N, is chosen to be 100 for the inverse design to ensure sufficient reflection of the rejected wavelengths. To guide the optimization, we define a loss function for gradient descent such that the transmission of each wavelength to the respective output Bragg filter is maximized and the reflection from the Bragg gratings back into the input port is minimized. This objective function is described in Equation 3. The variable is used to denote the transmission from the fundamental TE mode of the waveguide of wavelength from port to port where the input waveguide is port labeled port 0 and the two output ports are 1 and 2, the waveguides at the end of the Bragg filters. The coefficients and are used to weight the relative importance of minimizing reflection relative to maximizing the transmission through the device. The flow of optimization is as follows: first, the permittivity distribution of the 2D design region is kept continuous, then after the continuous structure is optimized, the structure is slowly binarized where each pixel (FDTD cell) is either silicon or oxide cladding. Lastly, feature size constraints are imposed in order to have a device that can be feasibly fabricated. For proof of concept, we design several WDM filters across the C- and L-bands of telecommunications, and their simulated transmission spectra are shown in Figure 2. This entire inverse design process is conducted using the Spins Photonics fdtdz software package.
3 Results
To analyze the benefit of the co-optimized inverse design approach, we examine the performance of inverse designed WDMs when the Bragg filters are not included in the optimization. The co-optimization case is compared with the traditional inverse design approach as well as when the Bragg filters are added to the structure after optimization. This analysis is detailed in Figure 3. A significant reduction in crosstalk and narrowing of bandwidth is observed for the co-optimized WDM and Bragg filter design as compared with simple inverse design (Figure 3d.). By including the Bragg filters in the large scale simulation region used for optimization, the resulting structure reduces the reflection back into the input port while also efficiently routing the wavelengths to their respective outputs (difference shown in Figure 3 b.,c., and e.). A true advantage to using Bragg filters is the ability semi-arbitrarily reduce crosstalk by simply increasing the number of Bragg periods. This can be done at no expense to insertion loss. This effect is explored in Figure 3a. where we measure the same devices but with increasing Bragg filter lengths (indicated by the number of Bragg periods N). By increasing N to 300, crosstalk of less than -40 dB can be achieved while maintaining insertion loss. This approach has a minimal impact on the overall device footprint, as it involves merely extending the already narrow output waveguides.
Additionally, we explore the application of the WDM devices at a system-level. One of the WDM devices is coupled to a silicon nitride frequency comb source and the comb lines are separated into each of the ports. The frequency comb source is generated using a photonic crystal microring resonator giving the input spectrum shown in Figure 4b. and the resulting output spectra from the two ports of the WDM are shown in Figure 4c. Four comb lines are coupled into the output of one port of the WDM while three comb lines are coupled into the other port. As observed in Figure 4c., the pump for the frequency comb is coupled to both of the WDM outputs due to the higher power relative to other comb lines.
4 Discussion
Through co-optimization of distributed Bragg gratings and dielectric wavelength division multiplexers, we achieve low insertion loss and narrowband operation with ultra-low crosstalk. This design process is compatible with foundry fabrication protocols and easily transferable to different materials and material stacks. Additionally, this approach can be scaled to more than two output ports and other wavelength regimes simply through the redesign of the Bragg filters. As optical transceiver research pushes data rates beyond 1 Tbps [19, 29], optical power may increase beyond the material limits of silicon [30]. Materials like silicon nitride are CMOS foundry compatible and have lower waveguide loss, two-photon absorption coefficient and are more thermally stable than silicon [31]. Yet, it is difficult to create narrow-band inverse design WDMs in silicon nitride due to the lower index contrast with oxide cladding as compared with silicon. Our design approach could be used to overcome this limitation for narrow-band WDMs in lower index materials. Our designs can be readily translated to foundry fabrication processes and incorporated in optical interconnects. With our approach, we demonstrate the highest crosstalk suppression and narrowest channel spacing of dielectric inverse design wavelength division multiplexers to date.
5 Methods
5.1 Device Fabrication
The designed devices are fabricated on a silicon on insulator (SOI) substrate with a silicon device layer of 220 nm. The inverse-designed WDM and Bragg structures are patterned using electron beam lithography and positive resist (ZEP520A). The resist is developed in solvents (xylene then methyl isobutyl ketone (MIBK) and isopropyl alcohol (IPA)) and the silicon is etched using reactive ion etching (RIE). A HBr and Cl2 chemistry is used for the RIE. The resist is removed (using Remover 1165) and the substrate is cleaned in piranha solution (sulfuric acid and hydrogen peroxide) and an HF dip prior to oxide cladding. The silicon structures are cladded in oxide through spin coating of hydrogen silsesquioxane (HSQ) and annealing in nitrogen gas at 900∘ C.
5.2 Experimental Setup
The transmission characteristics of the devices are measured experimentally using a fiber-to-chip grating coupler setup. Input light from a telecom laser (Toptica, 1510-1630 nm) is swept continuously while the output power from the WDM ports is measured simultaneously on an optical power meter (Newport 1919-R). The transmission of the devices is normalized to the output power spectrum of the two grating couplers, to abstract the performance of the WDM device alone.
6 Acknowledgements
6.1 Funding
This work is funded by DARPA under the PIPES program and AMD through the System-X Alliance at Stanford. J.G. acknowledges support from the Hertz Foundation Graduate Fellowship. S.E. acknowledges support from the Shoucheng Zhang Graduate Fellowship and the Korea Foundation for Advanced Studies Overseas Ph.D. Fellowship. Part of this work was performed at the Stanford Nanofabrication Facility (SNF) and the Stanford Nano Shared Facilities (SNSF), supported by the National Science Foundation under Grant No. ECCS-2026822.
6.2 Author Contributions
S.M., G.A., and J.V. conceived of the project. S.M. designed, simulated, and measured the devices. S.M. and S.E. fabricated the devices. J.G. designed and fabricated the frequency comb source, and J.G. and S.M. conducted the frequency comb measurements. G.A. and J.V. supervised the project. All authors contributed to the writing of the manuscript and the analysis of the results.